• Main INDEX
  • Monthly INDEX
  • NEXT

    User name M JOnes

    Log entry time 00:14:07 on June 1,2002

    Entry number 4316

    keyword=Swing shift summary

    4:00 No beam . MCC working.
    5:00 MCC is attempting to deliver 100uA.
    Failure to deliver 100uA and had to settle for 67uA
    18:11 run 1035 thin 12C sieve. Nice sieve pattern.
    The multifoil target shows the sieve pattern cutoff. Must be a problem
    with the target or maybe the beam is hitting the target low. Decide to do dummy target runs.
    19:25 run 1036 9 foil confirm that large positive theta angles were being blocked
    10:26 run 1037 4 cm dummy nice sieve !
    20:43 run 1038 15 cm dummy nice sieve !

    Change momentum to 2.45 Gev/c -2%
    21:11 run 1039 opt target cut off
    21:34 run 1040 thin carbon nice sieve!
    22:11 run 1041 4cm dummy nice sieve!

    Call Dave Meekins to sort out target problems. He goes into the hall and
    sees that it isn't a major failure of target motion.

    Can't do arc energy measurement because Field Measurement gives message
    PROBLEM ---> pb read status.